# a Benchmark Dataset for Diagnoses and Deterioration Prediction in Emergency Medicine

### v3

@@ 3.3 Limitations
Despite the promising results, there are several limitations to consider. Firstly, the presented models rely on the specifics of Beth Israel Deaconess Medical Center in terms of patient demographics, available resources, process characteristics, and data collection practices. We carefully quantify the statistical uncertainties of our results and are convinced of the robustness of our results for this particular scenario within the limits of a retrospective evaluation. However, we cannot make any statements about the transferability of our findings to other hospitals or deployment scenarios as this would require an external validation dataset, see [19] for a recent perspective on the role of external validation for clinical decision support. Secondly, the reliance on data collected within the first 1.5 hours of the ED visits may limit the model’s applicability. We deliberately decided to focus on this setting to keep the dataset homogeneous. However, there is also a clear clinical case for a prediction model only incorporating information available during triage. A similar dataset to the one proposed could be created, but this task is left for future research. Thirdly, while we focus on the construction of a benchmark dataset and provide the first results for it, we do not provide any insights into the trained models beyond a discussion of the model performance for specific conditions. This could be achieved for example through attribution methods proposed in the context of explainable AI, however, mostly in uni-modular settings [20], to verify that the models do not exploit spurious correlations to reach their strong performance. Again, a detailed investigation also of causal vs. associational attributions along the lines of [21] is beyond the scope of this submission. Finally, although the open-source dataset fosters innovation, the implementation of a large-scale single end-to-end model with a wide range of target variables such as the diagnoses case might result in challenges during deployment in resource-constrained hardware, see 9 in the supplementary material for a more extensive discussion.
@@ 3.4 Future research directions
The proposed dataset and benchmark protocol build on input parameters that are typically available in clinical workflows and address a set of comprehensive, clinically meaningful prediction tasks. We therefore envision it to represent a meaningful baseline to benchmark prediction algorithms in the field. An obvious extension would be to include free-text data along the lines of [16]. Going beyond MIMIC-IV, models developed on the proposed dataset could foreshadow the development of more comprehensive models on more specialized datasets such as [22], for example including molecular data as input and/or prediction targets, to address more fine-grained diagnostic or deterioration prediction targets.
@@ 4 Potential implications
5 Methods
5.1 Clinical workflow and dataset creation
Figure 1: Pipeline outlines the MDS-ED workflow, which involves feature collection encompassing patient demographics, biometrics (such as height, weight, and BMI), vital parameters and trends, laboratory values and trends, and ECG waveform data to address two clinically relevant prediction scenarios: predicting patient discharge diagnoses out of 1428 cardiac and non-cardiac ICD10-CM codes and predicting patient deterioration according to 15 clinical deterioration measures.
Figure 1 presents a schematic illustration of the proposed MDS-ED pipeline which task focuses on a specific workflow with high clinical relevance, where for each ED visit, it encompasses a set of features collected from a window of 90 minutes from the patient’s arrival at the ED to predict patient discharge diagnoses and deterioration through the stay. MDS-ED was created by linking ECG waveforms from the medical information mart for intensive care (MIMIC) database, specifically, MIMIC-IV-ECG [23] dataset to clinical features and outcomes as clinical ground truth from the MIMIC-IV and MIMIC-IV-ED dataset [24].
Prediction tasks and targets
In this work, we broadly categorize all prediction tasks into two groups: patient discharge diagnoses and patient deterioration. For discharge diagnoses, we follow MIMIC-IV-ECG-ICD [12], which proposed a framework to predict discharge diagnoses based on a single electrocardiogram (ECG). To this end, we frame the task as a multilabel classification task in terms of international classification of diseases clinical modification 10 (ICD-10 CM) codes. As in [12], we convert ICD-9 to ICD-10 CM codes where necessary, convert codes to 5 digits, and propagate codes to parent codes up to the third digit. In this way, we obtain a total set of 1428 significantly populated ICD-10 CM codes covering a wide range of cardiac and non-cardiac conditions. For patient deterioration, we build on a consensus definition [25] and investigate 15 deterioration events in a multilabel setting after the patient arrived at different intervals and after the 1.5-hour feature collection window: Six subtasks cover clinical deterioration within 24 hours and cover severe hypoxemia, extracorporeal membrane oxygenation (ECMO), use of vasopressors, use of inotropes, mechanical ventilation, and in-hospital cardiac arrest (IHCA). Two subtasks cover ICU admission within 24 hours and or within the entire stay, and finally, 7 subtasks cover mortality prediction at different horizons ranging from inpatient mortality, 24 hours, 7 days, 28 days, 90 days, 180 days, to 365 days. With these two main groups, we aim to cover a wide range spectrum of patient monitoring in the ED as well as long-term mortality across 1,443 unique target labels, see Table 5 in the supplementary material for a detailed target definition discussion.
Features
MDS-ED consists of data from 71,098 patients, encompassing 121,195 unique visits, and a total of 129,057 samples. We leverage the stratified splits provided along with the dataset accompanying [12], which includes stratification based on gender, age bins, and discharge diagnoses. We distribute the total number of 20 stratified folds into train, validation, and test following ratios of 18:1:1. For deep learning models with clinical routine data, we apply a median imputation learned from train set to validation and test for each column and we apply a masking strategy in which we add an additional binary column to indicate which values were imputed to the model. See the supplementary material for further preprocessing and definitions details as well as descriptive statistics on demographics, biometrics, vital parameters and laboratory values in Tables 6-8.
5.2 Related work
Group Detail MIMIC-ED-Assist VitalML MEME EHRSHOT MC-BEC MDS-ED
Source [13] [14] [18] [15] [17] This work
Population ED ED Longitudinal Longitudinal ED ED
Size Patients 25714 N/A 947028 6739 63389 71098
Visits 32356 19847 N/A 921499 102731 121195
Features Demographics
Biometrics
Vital parameters ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$)
Lab. values ✓($\mathbb{T}$) ✓($\mathbb{T}$)
Waveforms ✓($\mathbb{E}$)
Chief complaint
Medications
Tasks Labels 3 3 4 15 7 1443
Availability Open source
{tablenotes}
Table 4 contains a direct comparison with related works in terms of population, dataset size, features, availability, and number of target labels. [13] proposed the MIMIC-ED-Assist benchmark to reduce ED length of stay by flagging high-risk patients using triage features. [14] proposed VitalML to predict deterioration in the next 90 minutes from the first 15 minutes of monitoring. [18] proposed a multiple embedding model for electronic health records (EHR) (MEME) to generate pseudo-notes from tabular data and predict ED disposition, discharge location, ICU admission, and inpatient mortality utilizing a natural language processing approach. [15] proposed a few shot foundational model EHRSHOT approach to predict operational outcomes, developed diagnoses, as well as anticipate lab test results and image findings. [17] proposed MC-BEC to predict patient deterioration, disposition, and revisit from features collected including diagnoses from the first 15 minutes after the patient is assigned to a room.
5.3 Model architectures
To propose an initial benchmark, we experiment with diverse settings based on data modalities input such as (1) clinical routine data, (2) ECG features only, (3) ECG waveforms only, (4) ECG features + clinical routine data, and (4) ECG waveforms + clinical routine data. For waveforms (3) and (5), inspired by recent successful applications in the field of physiological time series [12, 27, 28] we employed an S4 classifier based on structured state space models with four layers [29] (with state dimension 8 and model dimension 512 as in [28]). For the multimodal models in (5), we use a 3-layer MLP as tabular encoder, where the three categorical features (gender, race, acuity score) are first processed through appropriate embedding layers. The outputs of the signal encoder and the tabular encoder are then fused through concatenation. For (1), (2) and (4) we fit extreme gradient boosting (XGBoost) decision trees. Additional details on model architectures and model hyperparameters can be found in 9 Table 9.
5.4 Training and evaluation
For the S4 models, we used Adam Weight Decay Optimization (AdamW) as the optimizer, with both learning rate and weight decay set to 0.001, and maintained a constant learning rate schedule. The training was conducted with a batch size of 64 samples over 20 epochs which usually converged earlier and model selection on the validation set. The training objective was to minimize binary cross-entropy loss. For all the models, we evaluate performances on the macro average across all areas under the respective receiver operating curves (AUROC) (macro AUROC). To assess statistical uncertainty resulting from the test set’s finite size and specific composition, we employ empirical bootstrap on the test set with $n=1000$ iterations. We report 95% confidence intervals for both macro AUROC and individual label AUROCs. Given [12] evaluation protocol, we use only the first record per visit per patient to avoid bias. Additional details of training and evaluation procedures can be found in 9.
6 Availability of source code and requirements
@@
  • Project code repository: [31].
@@
  • License: MIT
7 Declarations
7.1 List of abbreviations
AdamW: Adam Weight Decay Optimization, AI: Artificial Intelligence, AUROC: Area Under the Receiver Operating Characteristic Curve, ECG: Electrocardiogram, ECMO: Extracorporeal Membrane Oxygenation, ED: Emergency Department, EHR: Electronic Health Records, ICD CM: International Classification of Diseases Clinical Modification, ICU: Intensive Care Unit, IHCA: In-Hospital Cardiac Arrest, MDS-ED: Multimodal Decision Support for Emergency Department, MIMIC: Medical Information Mart for Intensive Care, MLP: multi-layer perceptron, XGBoost: Extreme Gradient Boosting.
7.2 Consent for publication
Not applicable
@@ 7.3 Competing Interests
7.4 Author’s Contributions
JMLA and NS conceived and designed the project. NS and JMLA implemented the prediction models. JMLA conducted the experimental analyses, with NS supervising them, and HB providing critical revision for clinical intellectual content. JMLA produced the first draft, NS and HB revised it. All authors critically revised the content and approved the final version for publication.

### v4

@@ 4.3 Limitations
Despite the promising results, there are several limitations to consider. Firstly, diagnostic labels were inferred from ICD10 codes, which are subject to biases and incompleteness as they are primarily used for billing purposes and are not assigned purely based on clinical considerations. Clinical deterioration labels are comparably reliable as they are sourced from the same hospital system. Post-discharge mortality might again be subject to inaccuracies as it originates from a different source system.
Secondly, the presented models rely on the specifics of Beth Israel Deaconess Medical Center in terms of patient demographics, available resources, process characteristics, and data collection practices. The presented results allow statements about the generalizability of the proposed models to unseen patients with well-defined statistical uncertainty measures, as far as this is possible. However, it does not allow any statements about the transferability of these findings to other hospitals. It is important to note that generalization to other sites is challenging not only due to different patient demographics but also due to available resources, process characteristics, and data collection practices. This extends as far as certain labels, such as ICU admission, that are directly affected by process characteristics. The challenges of external validation in the context of complex clinical decision support systems are nicely summarized in a recent article [30].
Thirdly, all considered prediction scenarios rely on data collected within the first 1.5 hours of the ED visits, which may limit the model’s applicability. We deliberately decided to focus on this setting to keep the dataset homogeneous. There are two directions that are worth exploring in future work. On the one hand, there is a clear clinical case for a prediction model only incorporating information available during triage. On the other hand, it might be instructive to extend the time horizon beyond 1.5 hours in particular to improve the performance for deterioration prediction tasks. In both cases, datasets similar to the one proposed could be created, but this task is left for future research.
Finally, structured benchmarks through open-source datasets is only a first step towards measurable progress in the domain. Translating these research findings into clinical applications poses significant additional challenges. For example, deploying large-scale models will require substantial investments in hardware, in particular for real-time applications. Clinical decision support systems have to be carefully validated on diverse patient population and have to undergo certification, have to be properly integrated into clinical workflows and have to be closely monitored after deployment to detect early signs of performance deterioration due to time-dependent dataset shifts.
To ensure clinical utility, future work should prioritize validation across diverse patient populations and healthcare settings. While many AI algorithms are developed and published, only a small fraction undergo prospective validation or certification for real-world deployment. Regulatory approval and integration into existing clinical workflows present significant challenges, which may impact clinician trust and adoption. A structured benchmarking process could help to streamline this by ensuring a more standardized performance evaluation.
@@ 4.4 Future research directions
The proposed dataset and benchmark protocol are based on input parameters that are typically available in clinical workflows and address a set of comprehensive and clinically meaningful prediction tasks. We therefore envision it to represent a meaningful baseline to benchmark prediction algorithms in the field. In addition to addressing the limitations mentioned in the previous sections, we consider pursuing the following technical research directions as particularly promising:
Future research directions could focus on expanding the scope of input variables beyond the current feature set, incorporating additional clinical and physiological signals to enhance predictive performance. An obvious extension along these lines would be to include free-text data as in the recent of Hager et al [17]. This could enhance the predictive performance by capturing richer clinical context, uncovering implicit relationships, and incorporating nuanced information that structured data alone might miss. Going beyond MIMIC-IV, models developed on the proposed dataset could foreshadow the development of more comprehensive models on more specialized datasets such as [31], for example including molecular data as input and/or prediction targets, to address more fine-grained diagnostic or deterioration prediction targets. Another promising avenue involves using multiple waveform snapshots over time for deterioration prediction, rather than relying solely on a single observation, thereby capturing dynamic changes in patient conditions more effectively.
Finally, our work naturally focuses on quantitative accuracy. However, there are many quality dimensions beyond quantitative accuracy that are important for clinical applications, such as explainability, different aspects of robustness, uncertainty quantification or privacy. As a specific example, it would be interesting to investigate whether specific ECG patterns contribute to the performance gains observed when using full ECG waveforms instead of extracted features. Related methods for explaining ECG-based predictions have been explored in [32], and further work could extend these approaches to assess causal versus associational attributions along the lines of [33]. As a second specific example, even though we work with de-identified structured data in this work, it is important to develop models in line with privacy requirements. Here, emerging privacy-preserving techniques, as detailed for example in [34], will become essential to ensure compliance in clinical decision support systems, especially with the potential future inclusion of free-text data.
@@ 5 Potential implications
5.1 Data and code availability
  • 1.
@@ 5.1 Data and code availability
  • 3.
    Project code repository: https://github.com/AI4HealthUOL/MDS-ED
  • 4.
@@
5.2 Author contributions
JMLA and NS were responsible for the conceptualization of the project. Data curation was performed by JMLA and NS. JMLA and NS implemented the prediction models, with JMLA conducting the formal analysis under NS’s supervision. HB provided critical revisions for clinical content. JMLA produced the original draft, and NS and HB contributed to the review and editing. All authors participated in the investigation, methodology, and validation, and they approved the final version for publication. Computing resources were provided through Carl von Ossietzky University of Oldenburg.
5.3 Generative AI statement
Generative AI was used for grammar correction but not for text generation in this manuscript.
@@ 5.4 Competing interests
5.5 Ethics statement
This study used publicly available and deidentified datasets. No direct patient interaction or intervention was involved. As these datasets are released under established data use agreements and have been ethically approved for secondary research, no additional ethics approval was sought.
Appendix A Effect of the binary NaN mask on model performance
For the ECG waveforms + clinical routine data setting (deep learning), our main results achieved 0.8256 (0.8222, 0.8288) for the diagnoses model and 0.9115 (0.8991, 0.9222) for the deterioration model when using a binary NaN mask column. In this section, we present additional experiments as a small ablation study to demonstrate that omitting the binary NaN mask column leads to inferior performance. Specifically, without the mask, the models achieved 0.8150 (0.8113, 0.8185) and 0.8986 (0.8864, 0.9094), respectively.
Appendix B Advantages of multimodal integration in clinical settings
In Table 5, we present a more finegrained performance comparison, which disentangles the overall performance according to ICD chapters in the diagnostic case and according to three categories, see Table 5 for the assignment of individual deterioration labels to these categories. In Table 6, we show a detailed breakdown of the model performance within the deterioration case. We focus on highlighting the added value of including full waveforms (ECG W. & CR) as opposed to waveform features (ECG F. & CR). We indicate the relative improvement of the former over the latter setting in the final column in both tables. Most notably, including full waveforms yields performance improvements for every single entry in both tables.
Table 5: Detailed breakdown of model performance for different ICD-10 categories and clinical deterioration outcomes. Unimodal settings are CR: features derived from clinical routine data, ECG F.: ECG features, and ECG W.: ECG waveforms. Multimodal settings are ECG F. & CR: ECG features and clinical routine data, as well as ECG W. & CR: ECG waveforms and clinical routine data. Finally, Relative Improvement (%): relative improvement of ECG W. & CR against ECG F. & CR.
CR ECG F. ECG W. ECG F. & CR ECG W. & CR Relative Improvement (%)
IX: Circulatory 0.7645 0.7153 0.8050 0.8050 0.8761 8.83
II: Neoplasms 0.8109 0.5886 0.8101 0.8101 0.8672 7.05
XIX: Injury, Poisoning 0.7425 0.6262 0.7712 0.7712 0.8479 9.95
I: Infections 0.7485 0.6258 0.7769 0.7769 0.8346 7.43
IV: Endocrine 0.7723 0.6401 0.7920 0.7920 0.8336 5.25
III: Blood 0.8048 0.5992 0.8143 0.8143 0.8335 2.36
XI: Digestive 0.7510 0.5902 0.7604 0.7604 0.8270 8.76
XIV: Genitourinary 0.7743 0.6053 0.7766 0.7766 0.8253 6.27
X: Respiratory 0.7300 0.6581 0.7643 0.7643 0.8226 7.63
V: Mental 0.7398 0.6456 0.7619 0.7619 0.8111 6.46
XII: Skin 0.6951 0.6284 0.7153 0.7153 0.8087 13.06
VI: Nervous 0.7088 0.5870 0.7183 0.7183 0.7955 10.75
XVIII: Symptoms 0.6983 0.6090 0.7168 0.7168 0.7813 9.00
XIII: Musculoskeletal 0.6786 0.5939 0.6930 0.6930 0.7405 6.85
Clinical Deterioration 0.7729 0.7135 0.7515 0.8294 0.9070 9.36
ICU Admission 0.8738 0.7377 0.7474 0.8865 0.9063 2.23
Mortality 0.8792 0.7579 0.8160 0.8920 0.9168 2.78
Table 6: Detailed breakdown of model performance category for all investigated sets of input features in the deterioration setting. Unimodal settings are CR: features derived from clinical routine data, ECG F.: ECG features, and ECG W.: ECG waveforms. Multimodal settings are ECG F. & CR: ECG features and clinical routine data, as well as ECG W. & CR: ECG waveforms and clinical routine data. Finally, Relative Improvement (%): relative improvement of ECG W. & CR against ECG F. & CR. We summarize the first six entries as ‘Clinical Deterioration’, the following two as ‘ICU Admission’ and the remaining seven as ‘Mortality’.
CR ECG F. ECG W. ECG F. & CR ECG W. & CR Relative Improvement (%)
Severe hypoxemia 0.4834 0.5311 0.7030 0.5055 0.6980 38.08
ECMO 0.7873 0.6126 0.7768 0.8991 0.9355 4.05
Vasopressors 0.8241 0.7759 0.7672 0.8551 0.9239 8.05
Inotropes 0.6550 0.7900 0.6184 0.8064 0.9400 16.57
Mechanical ventilation 0.9373 0.7262 0.7673 0.9418 0.9590 1.83
IHCA 0.9502 0.8454 0.8760 0.9682 0.9859 1.83
ICU 24 hours 0.8809 0.7390 0.7443 0.8928 0.9147 2.45
ICU overall 0.8666 0.7363 0.7506 0.8801 0.8979 2.02
In-hospital 0.8371 0.8753 0.8715 0.8736 0.9423 7.86
24 hours 0.9304 0.7267 0.8409 0.9304 0.9600 3.18
7 days 0.9176 0.7266 0.8176 0.9286 0.9429 1.54
28 days 0.8858 0.7329 0.7859 0.8946 0.9115 1.89
90 days 0.8651 0.7590 0.7926 0.8792 0.8952 1.82
180 days 0.8605 0.7464 0.8036 0.8727 0.8894 1.91
365 days 0.8579 0.7387 0.7999 0.8646 0.8768 1.41
Appendix C Dataset details
Table 7: Definitions of labels for the diagnoses and deterioration models, including 1428 discharge diagnoses based on ICD10-CM codes in the diagnoses model, and 15 labels covering clinical outcomes such as mortalities, ICU admission, vasopressors, inotropes, mechanical ventilation, in-hospital cardiac arrest (IHCA), extracorporeal membrane oxygenation (ECMO), and severe hypoxemia in the deterioration model.
Name Labels Details
Diagnoses model (1428 labels, based on [12])
Diagnoses 1428 Discharge diagnoses ICD10-CM codes
Deterioration model (15 labels, based on [23])
Mortalities 6 1, 7, 30, 90, 180, 365 days and in-stay
ICU Admission 2 24 hours and in-stay
Vasopressors 1 Within 24 hours: epinephrine, norepinephrine, vasopressin, dobutamine, dopamine, or phenylephrine
Inotropes 1 Within 24 hours: epinephrine, dobutamine, or dopamine
Mechanical Ventilation 1 Same day or next day ICD9/ICD10 codes: 9670, 9671, 9672, 5A1935Z, 5A1945Z, and 5A1955Z
In-Hospital Cardiac Arrest (IHCA) 1 Same day or next day ICD9/ICD10 codes: I469, 4275, I462, V1253, I468
Extracorporeal Membrane Oxygenation (ECMO) 1 Same day or next day ICD9/ICD10 codes: 3961, 3965, 3966, 5A1221Z, 5A1522G, 5A1522H, 5A15223, 5A1522F, 5A15A2F, 5A15A2G, and 5A15A2H
Severe Hypoxemia 1 Oxygen saturation $\leq$ 85% within 24 hours
Table 7 contains the considered tasks by description, number of labels, and additional details such as target criteria definition. Apart from diagnoses, we use the rest of the tasks defined as patient deterioration based on previous work from the literature that defines physiological deterioration [23]. Although we consider diverse scores such as SOFA significant, calculating these markers requires a more detailed definition of downstream task prediction, which goes beyond the scope of maintaining homogeneity across tasks in this study. The curation of the dataset starts from MIMIC-IV-ECG-ICD [12], from which we obtained the ICD10 diagnoses codes for the complete MIMIC-IV-ECG [20] dataset, therefore, based on ECG records with available ED stay_id, we select the ones that happened between the patient in time and the 1.5 hours window. Mortalities at finite horizons were computed based on the date of death from [20], and for the computation of mortality per stay, we include discharge date times from the admissions table in [21] MIMIC-IV to account for final discharge, all of the mortalities dates were previously validated from the source with a follow-up up to 1 year after the patient’s last recorded discharge. ICU admission in 24 hours and per stay was computed based on the ICU stays table from [21]. Mechanical ventilation and ECMO were computed given the procedures table from MIMIC-IV-ED given the patient admission and procedure date. Cardiac arrest was computed from the diagnosis tables from MIMIC-IV-ED and MIMIC-IV, where we first validated if the patient was discharged with the diagnosis code within the ED or hospital otherwise. We clarify that although mechanical ventilation, ECMO, and cardiac arrest do not provide an exact time of events, we carefully approximate the times with dates given the same day or the day after. Vasopressors and inotropes were computed from the pyxis table from MIMIC-IV-ED given the exact time of the applied medication. Finally, severe hypoxemia was computed from the vital signs table from MIMIC-IV-ED. In the deterioration setting, to not discard samples where the events happened before the 1.5 hours for a single task, e.g. severe hypoxemia within 1.5 hours, we include a special token for the task in the label space which we mask during training and evaluation in order not to negatively affect training and evaluation for other labels due to exclusion of entire patients, see D for more details.
Table 8: Summary demographic characteristics across samples including gender distribution, age distribution (median with standard deviation and quartiles), and ethnicity distribution in the study population, presented as percentages.
Name Values
Gender (%)
Female 68289 (52.9)
Male 60806 (47.1)
Age (%)
Median years (SD) 64 (28)
Quantile 1: 18-49 32634 (25.27)
Quantile 2: 50-64 34376 (26.62)
Quantile 3: 65-77 31821 (24.64)
Quantile 4: 78-101 30264 (23.44)
Ethnicity (%)
White 78310 (60.66)
Black 27680 (21.44)
Hispanic 9230 (7.14)
Asian 4700 (3.64)
Other 9175 (7.10)
Table 8 provides detailed demographic information on the study population, including gender, age, and ethnicity. The gender distribution shows a slightly higher proportion of females (52.9%) compared to males (47.1%). The age distribution is divided into four quantiles, with a median age of 64 years and a standard deviation of 28 years. Regarding ethnicity, the majority of the population is White (60.66%), followed by Black (21.44%), Hispanic (7.14%), Asian (3.64%), and individuals of other ethnicities (7.10%). This diverse demographic distribution ensures a comprehensive evaluation of the model across various patient groups. These variables were extracted from the ed stays table from MIMIC-IV-ED.
Table 9: Summary of statistics across samples for biometric measurements, vital parameters, and triage acuity levels, including minimum, maximum, median, interquartile range (IQR), and number of records for each variable.
Name Unit Minimum Maximum Median IQR Records
Biometrics
BMI kg/m2 0 99.2 27.7 8.70 71234
Weight kg 20.41 367.99 77.11 27.56 74026
Height cm 64.99 387.09 167.64 15.23 42084
Vital parameters
Temperature C 13.3 44.72 36.6 0.55 77179
Heart rate bpm 0.0 570.0 84.0 30.0 153249
Respiratory rate bpm 0.0 169.0 18.0 4.0 150525
Oxygen saturation % 0.0 100.0 99.0 3.0 146140
SBP mmHg 0.0 274.0 130.0 35.0 151346
DBP mmHg 0.0 486.0 73.0 23.0 151254
Triage
Acuity level 1 5 2 1 126071
Table 9 provides details on the biometrics, vital parameters, and triage values considered in this work, including the name, unit, minimum value, maximum value, median value, and the number of records. Dataset preprocessing includes weight and height conversion from pounds and inches to kilograms and centimeters. During the data cleaning process, we removed outliers and introduced missing values at various thresholds to ensure model computational convergence. The specific thresholds for outliers were as follows: temperature below and above 50 and 150 Fahrenheit degrees respectively. Heart rate above 700, respiration rate above 300, oxygen saturation below 0 and above 100, diastolic and systolic blood pressures above 500. BMI above 100, weight below 20 or above 400, and height below 60 and above 400. The temperature unit was converted from Fahrenheit to Celsius, whereas weight from pounds to kilograms, and height from inches to centimeters. Biometrics features were extracted from the online medical records table from MIMIC-IV from which we match the closest value of a patient stay within 30 days of admission, otherwise missing. Vital parameters and acuity were extracted from the vital signs and triage tables of MIMIC-IV-ED.
Table 10: Summary of laboratory values across samples considered in this study, including the name, unit, minimum value, maximum value, median value, fluid type, and the number of records for each parameter.
Name Unit Minimum Maximum Median IQR Fluid Records
Abs. Basophil Count K/uL 0.0 9.2 0.04 0.03 Blood 36837
Abs. Eosinophil Count K/uL 0.0 15.33 0.09 0.15 Blood 36838
Abs. Lymphocyte Count K/uL 0.0 88.08 1.47 1.12 Blood 36909
Alanine Aminotransf. IU/L 1.0 1999.0 22.0 22.0 Blood 29999
Albumin g/dL 1.0 6.7 4.0 0.79 Blood 28324
Alkaline Phosphatase IU/L 5.0 1985.0 85.0 51.0 Blood 29923
Asparate Aminotransf. IU/L 1.0 1996.0 29.0 29.0 Blood 29995
Bands % 0.0 61.0 0.0 3.0 Blood 6057
Base Excess mEq/L -413.0 29.0 0.0 6.0 Blood 8011
Basophils % 0.0 35.0 0.4 0.39 Blood 79386
Bicarbonate mEq/L 2.0 50.0 24.0 5.0 Blood 79771
Bilirubin, Direct mg/dL 0.0 45.5 1.0 2.5 Blood 1174
Bilirubin, Total mg/dL 0.0 55.8 0.5 0.5 Blood 30098
C-Reactive Protein mg/L 0.1 299.3 17.95 75.14 Blood 1108
Calcium, Total mg/dL 0.0 79.2 9.2 0.79 Blood 30262
Carboxyhemoglobin % 0.0 18.0 2.0 2.0 Blood 1078
Chloride mEq/L 58.0 140.0 102.0 6.0 Blood 79855
Creatine Kinase (CK) IU/L 8.0 1998.0 115.0 158.0 Blood 9588
CK, MB Isoenzyme ng/mL 1.0 497.0 3.0 4.0 Blood 9272
Creatinine mg/dL 0.0 43.0 1.0 0.5 Blood 82850
Eosinophils % 0.0 75.0 1.3 2.1 Blood 79387
Fibrinogen, Functional mg/dL 35.0 1348.0 298.5 161.5 Blood 2280
Free Calcium mmol/L 0.48 3.31 1.08 0.12 Blood 1472
Glucose mg/dL 3.8 1684.0 113.0 50.0 Blood 85022
Hematocrit % 6.9 66.1 38.3 7.89 Blood 82490
Hemoglobin g/dL 0.0 21.8 12.6 3.0 Blood 84890
INR(PT) nan 0.75 24.0 1.1 0.30 Blood 44656
Lactate mmol/L 0.3 25.2 1.8 1.20 Blood 32666
Lymphocytes % 0.0 100.0 19.7 17.1 Blood 79387
Magnesium mg/dL 0.2 45.0 2.0 0.40 Blood 30472
Neutrophils % 0.0 99.0 69.9 19.59 Blood 79387
Oxygen Saturation % 0.2 901.0 80.0 32.0 Blood 4838
PT sec 8.2 150.0 12.0 3.5 Blood 44656
PTT sec 16.6 150.0 30.6 6.99 Blood 44497
Phosphate mg/dL 0.2 23.8 3.4 1.1 Blood 29888
Platelet Count K/uL 5.0 1898.0 227.0 105.0 Blood 82260
Potassium mEq/L 1.3 10.7 4.2 0.80 Blood 79857
RDW % 10.6 33.8 13.8 2.19 Blood 82266
Red Blood Cells m/uL 0.65 8.06 4.25 0.96 Blood 82266
Sodium mEq/L 88.0 179.0 139.0 5.0 Blood 79865
Troponin T ng/mL 0.0 17.36 0.06 0.13 Blood 41806
Urea Nitrogen mg/dL 1.0 263.0 18.0 13.0 Blood 82522
White Blood Cells K/uL 0.0 632.1 8.1 4.60 Blood 82367
pCO2 mm Hg 8.0 189.0 44.0 17.0 Blood 8008
pH units 5.0 9.0 6.0 1.0 Urine 20968
Table 10 provides details on the laboratory values considered in this work, including the name, unit, minimum value, maximum value, median value, fluid type, and the number of records. During the data cleaning process, we removed outliers and introduced missing values at various thresholds to ensure model computational convergence. The specific thresholds for outliers were as follows: absolute basophil count above 20, absolute eosinophil count above 20, absolute lymphocyte count above 100, alanine aminotransferase above 2000, alkaline phosphatase above 2000, aspartate aminotransferase above 2000, creatine kinase above 2000, glucose above 2000, lactate above 2000, and platelet count above 2000. Laboratory values were extracted from the laboratory events table from MIMIC-IV, where for each record we matched all the patient records of the desired tests available between patient record in-time and the 1.5 hours window.
Appendix D Models
Table 11: S4 model hyperparameters
Hyperparameter Value
Block of layers 4
S4 model copies 512
S4 state size 8
Optimizer Adam
Learning rate 0.001
Weight decay 0.001
Learning rate schedule constant
Batch size 64
Epochs 20
Table 11 outlines the hyperparameters employed in the S4 model. The architecture consists of four blocks of layers, with each block containing 512 copies of the S4 model. The state size within the S4 model is set to 8. For optimization, the AdamW optimizer is utilized with a learning rate and weight decay both set to 0.001. The learning rate schedule is maintained constant throughout training. A batch size of 64 samples is used for each training iteration, spanning a total of 20 epochs. The training objective is to minimize the binary cross-entropy loss. During training, we apply a model selection strategy on the best performance (AUROC) on the validation set which usually converges earlier than the final number of epochs. As previously described in C, with the goal of not discarding samples whose any of the deterioration targets happens before the 1.5-hour window or the target event cannot be clearly defined, we introduce a special token in the label space for which we account during and training and evaluation to be discarded. As of XGboost models, we fit these models with the default hyperparameter values in a binary setting for the later AUROC aggregation across settings, models which in principle do not intrinsically allow the inclusion of samples during training with the special token to be discarded.